A baseball is dropped from the roof of a building. After some time, it flies past a window of height 2.20m. if the time it takes the baseball to fall past the window is 1/3s, how far above the window is the roof?
1
Expert's answer
2019-08-29T09:52:36-0400

Let "v" be the velocity of ball just before it comes to the upper part of the window

Using 2nd equation of motion when the ball passing through window

"s=(u\times t)+(\frac{a}2\times t^2)"

now

"s=2.20m"

t"=\frac{1}{3}sec"

"a=g=9.8\frac{m}{{sec}^2}"


"2.20=(v\times{\frac1 {3}})+(\frac{9.8}{2}\times\frac{1}{3}\times\frac{1}{3})"


hence, "v=4.95\frac{m}{sec}"

Using 3rd equation of motion from starting point to window

"v^2-u^2=2\times a\times s"

"u=0" as ball started from rest

"v=4.95\frac{m}{sec}"

"a=g=9.8\frac{m}{{sec}^2}"

so,

"4.95^2=2\times 9.8\times s"

"s=1.25m"

Roof is "1.25m"  above the window